# Asymmetric Errors

### v1

@@ 4.1.4 Strength of a known mass peak: models of a known likelihood
A common analysis method is to take measurements of a reconstructed particle mass, and fit them to a total of signal and background shapes where all parameters are known except for the normalisations. Figure 15 shows such a situation, where the background is known to be flat and the signal is a Gaussian of mean 5.0 and sigma 1.0. One may then, as appropriate, work with the raw (unbinned) likelihood $\ln L=\sum_{i}\ln{(N_{B}B(x_{i})+N_{S}S(x_{i}))}$, where the sum is over all events, or the binned likelihood $\ln L=\sum_{j}F_{j}\ln n_{j}-F_{j}$, where $n_{j}$ is the contents of bin $j$ and $F_{j}$ is the prediction. The likelihood curves are shown in the second and third panels. In this example the unbinned likelihood gives a result $N_{S}={21.10}^{+9.39}_{-9.13}$, and the binned likelihood gives $N_{S}={19.54}^{+9.73}_{-9.46}$.
Figure 15: Fitting a peak with known shapes. The data (first panel) and the likelihoods, unbinned and binned, as a function of supposed signal strength
@@
Figure 16: Modelling the likelihoods in Figure 15.
@@ 4.1.5 A product with asymmetric uncertainties
The expected number of particle decays in some channel can be written as $N={\cal L}\sigma F$, where ${\cal L}$ is the integrated luminosity (or equivalent), $\sigma$ is the cross section [^1] It is regrettable to use the same symbol for cross-sections and for standard deviations, but the use is entrenched., and $F$ is the branching fraction. ($N$ is the ideal number: the actual number observed will be Poisson-distributed with mean $N$.) In a typical instance ${\cal L}$ will be accurately known, but $\sigma$ and $F$ may have asymmetric errors. What should we quote as the error for the expected $N$ given, say, ${\cal L}=1000{\rm pb}^{-1},\sigma={12.3}^{+.4}_{-.5}{\rm pb}$, and $F={0.12}^{+0.01}_{-0.02}$?
[…]
@@
4.2 Case Studies
4.2.1 Systematic uncertainties from many sources
LHCb has performed a Dalitz plot fit to decays of the $\Lambda^{0}_{b}$ [LHCbLambdab] looking at 22 contributions: we consider that of the $\Lambda(1800)$ as an illustration. Systematic uncertainties in the value arise from many sources, and are evaluated by toy Monte Carlo, varying the source and examining the spread of the result. They are listed in Table 7. The combined error using the dimidiated model, is ${\sigma^{+}}=0.05965,{\sigma^{-}}=0.03294$. Using the distorted model it is ${\sigma^{+}}=0.06098,{\sigma^{-}}=0.03485$. As one would expect, the largest contribution dominates the total.
FSource ${\sigma^{+}}$ ${\sigma^{-}}$
fix res +0.059 -0.029
amp model +0.001 -0.008
res +0.008 -0.015
finite acc +0.003 -0.003
acc model +0.001 -0.001
kin +0.001 -0.001
sWt pg +0.006 0.0
massfit comb +0.004 0.0
plus 6 other sources that are evaluated as zero
Table 7: Sources of systematic uncertainty, taken from [LHCbLambdab].
5 PDF or likelihood? Variances or limits?
As mentioned in Section 1.2, a pdf can give 68% central limits, but they are the horizontal lines on the Neyman confidence belt, not the vertical ones, and the relationship is subtle. For instance, if a pdf has a positive skewness so that some $a$ may give an $x$ with a large upward fluctuation, then the likelihood has a negative skewness as some $x$ may have come from a large upward fluctuation.
A common example occurs in the energy measurements from a calorimeter which have negative skew. A particle of energy $E$ has a good chance of recording an energy close to $E$, with fluctuations driven by photon statistics, but also a smaller chance of recording an energy significantly below $E$ due to losses (due to cracks, escapes, and dead regions); there are no corresponding mechanisms for increases in the recorded energy. For such a calorimeter the pdf for the recorded energy of a 10 GeV photon might typically have a peak at 10 GeV and 68% confidence region limits of $[8.0,10.2]GeV$. All straightforward. But the likelihood for the true energy of a particle with recorded energy 10 GeV has a positive skew: it has a good chance of being close to 10 GeV, a smaller chance of being somewhat above 10 GeV and having lost energy to a crack, and a negligibly small chance of coming much below 10 GeV.
Essentially the same problem occurs in the evaluation of confidence regions using the bootstrap. This technique can be used to approximate the pdf by taking repeated samples from the data, and these can be used to give a confidence region for some quantity. However this is not the confidence region for the true quantity, and to use it as such “amounts to looking up the wrong tables backwards”, as Peter Hall puts it [Edgeworth, p36] and [hall, p928].
A helpful example is the proportional Gaussian: $p(x)={1\over\sqrt{2\pi}\alpha a}e^{-{1\over 2}\left({x-a\over\alpha a}\right)^{2}}$. If a measurement is quoted as being accurate to 10% then a value of 100.0 could have arisen from a true value of 90.0 with a discrepancy of just over one sigma, or from a true value of 110 with a discrepancy of just under one sigma. The 68% central confidence region is $[90.9,111.1]$. For a given $a$ this is symmetric in $x$, but for a given $x$ it is not symmetric in $a$.
The Poisson is a a well-known example of a skew distribution, but it is unfortunately unhelpful. The probability distribution has a positive skewness (the outcome can fluctuate a long way upwards but downward fluctuations are limited by zero). The likelihood also has a positive skewness. But this is driven by the increase in variance with the mean, not by the skewness of the probability distribution (which works in the opposite direction).
5.1 What happens if you use the ‘wrong’ formalism.
@@ 5 PDF or likelihood? Variances or limits?
As as example we consider the combination of errors from two results, both with error ${\sigma^{+}}=2,{\sigma^{-}}=1$. These are shown in Table LABEL:tab:mixedup for several models. Not surprisingly – this is a large asymmetry – the different models give a spread of different values for the combined ${\sigma^{+}}$, and likewise for ${\sigma^{-}}$. But the pdf models are clustered separately from the $\ln L$ models. The values of ${\sigma^{+}}$ and ${\sigma^{-}}$ for the combination under the assumption that these come from pdfs are both larger than those under the assumption that these come from likelihoods. This discrepancy is not enormous, but it is larger than the differences between individual models. (The difference between any of these and the ’wrong’ method of summing separately in quadrature is somewhat larger.) The conclusion to be tentatively drawn is that if a pdf error is (wrongly) treated as a likelihood error, or vice versa, this makes a meaningful, though not enormous, difference.
Similarly Table LABEL:tab:mixedup2 shows the outcomes from combining two results: ${1.0}^{+2.0}_{-1.0}$ and ${2.0}^{+2.0}_{-1.0}$. Again, the result to be quoted depends on the choice of model, which is arbitrary (though the practitioner may choose to reject some models as being inappropriate for their use) so there is a spread. There are also overall shifts between the pdf treatment and the $\ln L$ treatment, which is not arbitrary. If you use the wrong treatment this will give wrong results, by an amount which is somewhat larger than the inescapable differences owing to the choice of model.
@@ 5.2 A final result
A typical final result is presented in the form $x^{+\sigma^{+}_{1}}_{-\sigma^{-}_{1}}{\ }^{+\sigma^{+}_{2}}_{-\sigma^{-}_{2}}$ where “the first error is statistical and the second is systematic”. The first is thus, presumably, an error obtained from the likelihood to some final fit, and the second from some pdf-based uncertainty obtained from combining many separate sources. In such cases one may still want to present a combined result: ${x}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$
As a pedagogical example we consider the determination of the total strength (in Becquerel, or counts per second) of an isotropic point source using a counter of intrinsic efficiency $\eta$. If $n$ counts are measured in time $t$ then the strength $R$ is obtained from $n$ through a conversion factor $A$.
$R=An={{4\pi\over S}{1\over\eta t}n}$ (15)
where $S$ is the solid angle presented by the counter to the source. We can suppose that $\eta$ and $t$ are known to high accuracy, but there is an error in $S$ due to an uncertainty in the distance between the source and the detector, measured as $x\pm\sigma$. This has been subjected to a standard OPAT analysis using a Monte Carlo simulation which accurately models the geometry. However it so happens that the detector consists of a circular disc of radius $r$, so the solid angle is $2\pi(1-{x\over\sqrt{r^{2}+x^{2}}})$.
In this example, if an efficient detector with $r=1$ cm at $x=5.0\pm 0.3$ cm from the source measures 50 counts in 1 hour, the algebra gives $A={103.0}^{+12.4}_{-11.6}$ and the result would be quoted as $R={5149.5}^{+763.0}_{-694.3}(stat.){}^{+618.1}_{-582.1}(syst.)$, the statistical error being the Poisson error on $n=50$, scaled by $A$, and the systemartic error the above error on $A$, scaled by $n$. The question then is how to combine these two errors.
Figure 17 shows the joint Pdf/likelihood function for the factor $A$. The asymmetric error quoted above is found from the Pdf at $x_{true}=5.0$. If instead we consider the likelihood at $x_{meas}=5$ (i.e. the vertical line), the $\Delta\ln L=-{1\over 2}$ prescription gives ${}^{+582.1}_{-618.0}$. This is close to (but not exactly) a simple interchange between ${\sigma^{+}}$ and ${\sigma^{-}}$. Going from the horizontal to the vertical on the confidence-belt inevitably involves this interchange: it may also involve other behaviour depending on the nature of the joint function.
If we (wrongly) treat the second, pdf-based, error as if it were a lnL error then the linear sigma model gives a total error of ${}^{+971.1}_{-915.6}$. If we do this but interchange ${\sigma^{+}}$ and ${\sigma^{-}}$ for the ‘systematic’ term we get ${}^{+957.8}_{-931.4}$. This is a ‘better’ representation of the total error, if one has to be given. The coverage for the central region should be 68.27%. Numerical studies of this case show the first ${}^{+971.1}_{-915.6}$ region has a coverage of 67.86 %, whereas the second ${}^{+957.8}_{-931.4}$ region has a coverage of 67.99 %, which is a (small) improvement.
Figure 17: Pdf/likelihood for the factor $A$, as a function of the measured value (horizontally) and the true value (vertically)
@@ 6 Conclusions
We trust we have convinced any reader of this paper that asymmetric errors are not to be taken up lightly: if differences appear between the ${\sigma^{+}}$ and ${\sigma^{-}}$ values in some part of an analysis, they will introduce considerable complication both technically and conceptually. If they are small and can legitimately be ignored then our advice would be to do so, rather than to retain them out of misplaced diligence.
If they cannot be avoided then they can be handled by the methods given. You need to be clear whether they are likelihood errors or pdf errors, and whether you are combining results or combining errors. You also need to choose models from among those given here, always taking more than one model to check the robustness of the result. Experience should show which model(s) are appropriate for a particular type of problem.
@@ 7 Acknowledgements
We would like to thank our colleagues for many discussions and suggestions, especially the PHYSTAT organisers: Lydia Brenner, Nick Wardle, Olaf Behnke, Sara Algeri and Louis Lyons (who suggested some of the examples). This paper arose out of a workshop at the Banff Center for Arts and Creativity in Alberta, Canada, and we are very happy to thank the staff for their organisation and hospitality. We are grateful to Anja Beck (and the LHCb collaboration) for the data used in Section 4.2.1
@@ Appendix A Modelling pdf errors
We present the details of possible models for near-Gaussian pdf functions. The list is not exhaustive, and others, such as the Generalised Extreme Value distribution, can also be suggested [Possolo]. Their motivations, and their technical challenges, vary widely. There is no universal ‘best’ model and the choice is ultimately down to the user. Use of more than one model in any application is strongly recommended as the differing results will give a feeling for the robustness of the whole procedure.
@@ A.1 The Dimidiated Gaussian
Figure 18: The Dimidiated Gaussian. The left hand plot shows the results of a simple OPAT analysis and the proposed approximation. The right hand plot shows the pdf projected on the vertical axis that would come from a Gaussian on the horizontal axis.
@@
$R=\begin{cases}M+{\sigma^{+}}\nu,&\nu\geq 0\\ M+{\sigma^{-}}\nu,&\nu\leq 0\\ \end{cases}$ (16)
The pdf for the variable $\nu$ is a unit Gaussian, so for $R$ the pdf is a dimidiated (or split) Gaussian with two equal halves, as shown on the right. This is sometimes referred to as a bifurcated Gaussian, but that is certainly inaccurate. The dimidiated Gaussian for $R$ is the probability transform of the unit Gaussian for $\nu$, under the model that they are related by Equation (16)
@@
$p(x)=\begin{cases}{1\over{\sigma^{-}}\sqrt{2\pi}}e^{-{1\over 2}\left({x-M\over{\sigma^{-}}}\right)^{2}}&\text{for }x<M\\ {1\over{\sigma^{+}}\sqrt{2\pi}}e^{-{1\over 2}\left({x-M\over{\sigma^{+}}}\right)^{2}}&\text{for }x>M\\ \end{cases}$ (17)
@@
$D=$ ${2\over 3S}\left(\sqrt{2\pi}\gamma-D^{3}({1\over\pi}-1)\right)$
@@
There is a limit to the asymmetry that can be accommodated by this model. If one of the Gaussians has zero width, then
@@
but such cases are extreme.

### v2

@@ 4.1.4 Strength of a known mass peak: models of a known likelihood
A common analysis method is to take measurements of a reconstructed particle mass, and fit them to a total of signal and background shapes where all parameters are known except for the normalisations. Figure 16 shows such a situation, where the background is known to be flat and the signal is a Gaussian of mean 5.0 and standard deviation 1.0. One may then, as appropriate, work with the raw (unbinned) likelihood $\ln L=\sum_{i}\ln{(N_{B}B(x_{i})+N_{S}S(x_{i}))}$, where the sum is over all events, or the binned likelihood $\ln L=\sum_{j}(F_{j}\ln n_{j}-F_{j})$where $n_{j}$ is the contents of bin $j$ and $F_{j}$ is the prediction. The likelihood curves are shown in the second and third panels of Figure 16. In this example the unbinned likelihood gives a result $N_{S}={39.41}^{+10.20}_{-10.13}$, and the binned likelihood gives $N_{S}={40.43}^{+10.40}_{-10.35}$.
Figure 16: Fitting a peak with known shapes: the data (first panel) and the likelihoods (second and third panels), unbinned and binned, as a function of supposed signal strength, showing the peak and the $\Delta\ln L=-{1\over 2}$ errors.
@@
Figure 17: Modelling the likelihoods in the case shown in Figure 16, plotting the differences from the true curve.
@@ 4.1.5 A product with asymmetric uncertainties
The expected number of particle decays in some channel can be written as $N={\cal L}\sigma F$, where ${\cal L}$ is the integrated luminosity (or equivalent), $\sigma$ is the cross section [^2] It is regrettable to use the same symbol for cross-sections and for standard deviations, but the use is entrenched., and $F$ is the branching fraction. $N$ is the ideal number and the actual number observed will be Poisson-distributed with mean $N$. In a typical instance ${\cal L}$ will be accurately known, but $\sigma$ and $F$ may have asymmetric errors. What should we quote as the error for the expected $N$ given, say, ${\cal L}=1000{\rm pb}^{-1},\sigma={12.3}^{+0.4}_{-0.5}{\rm pb}$, and $F={0.12}^{+0.01}_{-0.02}$?
The central value is $1000\times 12.3\times 0.12=1476$ events. Writing $\sigma=\sigma_{0}+\Delta_{\sigma},F=F_{0}+\Delta_{F}$ then, to leading order in a Taylor expansion, $N=N_{0}+({\cal L}F_{0})\Delta_{\sigma}+({\cal L}\sigma_{0})\Delta_{F}$, and the appropriate function (pdf or $\ln L$) is found by combining the errors, scaled by ${\cal L}F_{0}$ and ${\cal L}\sigma_{0}$ respectively. For the linear sigma model this gives ${1476}^{+136}_{-250}$. Using the linear variance model it is ${1476}^{+137}_{-251}$.
@@
5 Likelihood or pdf? Variances or Confidence?
5.1 Framework
An error quoted on a measurement may denote the rms of a pdf or a 68% central confidence region obtained from a likelihood. Previous sections have discussed the appropriate handling of these two separate cases. Now we consider the consequences if they are, inadvertently or deliberately, confused.
As mentioned in Section 1.2, a pdf can give 68% central limits, but they relate tothe horizontal lines on the Neyman confidence belt, not the vertical ones, and the relationship is subtle. In the example discussed in Figure 2 a pdf with positive skewness leads to a likelihood with negative asymmetry.Essentially the same problem occurs in the evaluation of confidence regions using the bootstrap. A pdf is approximatedby taking repeated samples from the data, and these can be used to give a confidence region for some quantity. However this is not the confidence region for the true quantity, and to use it as such “amounts to looking up the wrong tables backwards”, as Peter Hall puts it [19, p36] and [20, p928]. The proportional Gaussian: $p(x)={1\over\sqrt{2\pi}\alpha a}e^{-{1\over 2}\left({x-a\over\alpha a}\right)^{2}}$ has a symmetric pdf but a likelihood with positive asymmetry.The Poisson has a positive asymmetry for both probability and likelihood, as its variance increases with the mean. So there is no easy equivalence between the 68% confidence region obtained from the pdf and that from the likelihood.
5.2 What happens if you use the wrong formalism.
@@ 5.1 Framework
As an example we consider the combination of errors from two results, both with error ${\sigma^{+}}=2,{\sigma^{-}}=1$. These are shown in Table 11 for several models. Not surprisingly – this is a large asymmetry – the different models give a spread of different values for the combined ${\sigma^{+}}$, and likewise for ${\sigma^{-}}$. But the pdf models are clustered separately from the $\ln L$ models. The values of ${\sigma^{+}}$ and ${\sigma^{-}}$ for the combination under the assumption that these come from pdfs are both larger than those under the assumption that these come from likelihoods. This discrepancy is not enormous, but it is larger than the differences between individual models. (The difference between any of these and the ‘wrong’ method of summing separately in quadrature is somewhat larger.) The conclusion to be tentatively drawn is that if a pdf error is (wrongly) treated as a likelihood error, or vice versa, this makes a meaningful, though not enormous, difference.
${\sigma^{+}}$ ${\sigma^{-}}$
Pdf models
dimidiated 2.636 1.651
distorted 2.727 1.760
edgeworth 2.656 1.586
railway 2.715 1.775
$lnL$ models
linearsigma 2.467 1.526
linearvariance 2.562 1.562
cubic 2.000 1.464
logarithmic 2.530 1.550
constrainedquartic 2.363 1.492
Wrong 2.828 1.414
Table 11: Different outcomes from combining errors ${}^{+2}_{-1}$ and ${}^{+2}_{-1}$.
Similarly Table 12 shows the outcomes from combining two results: ${1.0}^{+2.0}_{-1.0}$ and ${2.0}^{+2.0}_{-1.0}$. Again, the result to be quoted depends on the choice of model, which is arbitrary (though the practitioner may choose to reject some models as being inappropriate for their use) so there is a spread. There are also overall shifts between the pdf treatment and the $\ln L$ treatment, which is not arbitrary. If you use the wrong treatment this will give wrong results, by an amount which is somewhat larger than the inescapable differences owing to the choice of model.
Result ${\sigma^{+}}$ ${\sigma^{-}}$
$\ln L$ models
azzalini 1.703 1.232 0.796
brokenparabola 1.800 1.166 0.892
constrainedquartic 1.676 1.239 0.784
cubic 1.634 1.193 0.771
linearsigma 1.673 1.244 0.760
linearvariance 1.668 1.255 0.738
logarithmic 1.670 1.251 0.745
matchedquintic 1.655 1.232 0.809
moldeddoublequartic 1.729 1.217 0.808
moldeddoublequintic 1.767 1.188 0.856
moldedquartic 1.729 1.234 0.809
PDG 1.673 1.244 0.791
poisson 1.661 1.262 0.720
seventh 1.730 1.226 0.881
simpledoublequartic 1.715 1.222 0.822
simpledoublequintic 1.775 1.186 0.878
Pdf models
dimidiated 1.703 1.318 0.825
distorted 1.758 1.363 0.880
edgeworth 1.732 1.328 0.793
railway 1.744 1.357 0.888
Table 12: Different outcomes from combining results ${1.0}^{+2.0}_{-1.0}$ and ${2.0}^{+2.0}_{-1.0}$.
@@ 5.3 A final result
A typical final result is presented in the form $x^{+\sigma^{+}_{1}}_{-\sigma^{-}_{1}}{\ }^{+\sigma^{+}_{2}}_{-\sigma^{-}_{2}}$ where “the first error is statistical and the second is systematic”. The first is thus, presumably, an error obtained from the likelihood to some final fit, and the second from some pdf-based uncertainty obtained from combining many separate sources. In such cases one may still want to present a combined result: ${x}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$.
As a pedagogical example we consider the determination of the total strength (in Becquerel, or counts per second) of an isotropic point source using a counter of intrinsic efficiency $\eta$. If $n$ counts are measured in time $t$ then the strength $R$ is obtained from $n$ through a conversion factor $A$
$R=An={{4\pi\over S}{1\over\eta t}n},$ (14)
where $S$ is the solid angle presented by the counter to the source.
We suppose that $\eta$ and $S$ are well known, but that there is a Gaussian error $\sigma_{t}=0.1s$ on the calibration of the time measurement $t$ which is systematic, i.e. described by a pdf, and gives an asymmetric error on $1/t$.
Considering an example where $A=100,\eta=1,n=50$ and $t=1s$, we would write $R={5000}^{+741}_{-674}{}^{+556}_{-455}$ where the first, ’statistical’, error comes from the Poisson errors on $n=50$ and the second ’systematic’ from the error on $1/t$.
If we, wrongly, treat the pdf error as if it were a likelihood error and combine the two errors this gives a combined result of ${5000}^{+906}_{-828}$ (using the linear-sigma model: other models give near-identical values). If we, wrongly, treat the lnL error as if it were a pdf and combine them, this gives ${5000}^{+909}_{-831}$ (using the dimidiated model: other models give near-identical values. There is little difference between the two outcomes.
As a validation we performed 10,000,000 toy experiments, generating values of $n$ from a Poisson distribution of mean 50 and values of $t$ from a Gaussian distribution of mean 1 and standard deviation 0.1, and applying the two prescriptions. The pdf method gave a coverage of $69.074\pm 0.015$%, only slightly above the nominal 68.269%, and the lnL method was even closer with a coverage of $68.387\pm{0.015}$%
It thus appears justifiable to combine (with caution) asymmetric ‘systematic’ and ‘statistical’ errors on a final result, treating them both either as one or the other.
@@ 6 Conclusions
We trust we have convinced any reader of this paper that asymmetric errors are not to be taken up lightly: if differences appear between the ${\sigma^{+}}$ and ${\sigma^{-}}$ values in some part of an analysis, they will introduce considerable complication both technically and conceptually. If these differencesare small and can legitimately be ignored then our advice would be to do so, rather than to retain them out of misplaced diligence. If the differences are large then the best course would be to report the full likelihood, at least for the parameter(s) of interest, or perhaps to provide future meta-analysts with some fuller description of the behaviour of the likelihood function.
If asymmetric errorscannot be avoided then they can be handled by the methods given. You need to be clear whether they are likelihood errors or pdf errors, and whether you are combining results or combining errors. You also need to choose models from among those given here, or ones of your own,always taking more than one model to check the robustness of the result. Experience should show which model(s) are appropriate for a particular type of problem.
This treatment assumes that the errors being considered are independent, and that the final result is a single quantity. We hope to consider its extension to correlated inputs and outputs in a future publication.
@@ 7 Acknowledgements
We would like to thank our colleagues for many discussions and suggestions, (including a PHYSTAT informal review), especially the PHYSTAT organisers: Lydia Brenner, Nick Wardle, Olaf Behnke, Sara Algeri and Louis Lyons (who suggested some of the examples). This paper arose out of a workshop at the Banff Center for Arts and Creativity in Alberta, Canada, and we are very happy to thank the Banff International Research Station (BIRS) and itsstaff for their organisation and hospitality. We are grateful to Hans Dembinski for helpful comments.
ARB acknowledges that the publication was produced with funding from the Italian Ministry of University and Research under the Call for Proposals related to the scrolling of the final rankings of the PRIN 2022 call – Project title “Latent variable models and dimensionality reduction methods for complex data” – Project No. 20224CRB9E, CUP C53C24000730006 – PI Prof. Paolo Giordani – Grant Assignment Decree No. 1401 adopted on 18.9.2024 by MUR.
@@ Appendix A Modelling pdf errors
We present the details of possible models for near-Gaussian pdf functions. The list is not exhaustive, and others, such as the Generalised Extreme Value distribution, can also be suggested [10]. Their motivations, and their technical challenges, vary widely. There is no universal ‘best’ model and the choice is ultimately down to the user. Use of more than one model in any application is strongly recommended as the differing results will give a feeling for the robustness of the whole procedure.
@@ A.1 The dimidiated Gaussian
Figure 18: The Dimidiated Gaussian. The left hand plot shows the results of a simple OPAT analysis that gave a result ${25.0}^{+2.0}_{-1.5}$,and the proposed approximation. The right hand plot shows the resultingpdf projected on the vertical axis that would come from a Gaussian distribution for $\nu$ on the horizontal axis.
@@
$R=\begin{cases}M+{\sigma^{+}}\nu,&\nu\geq 0\\ M+{\sigma^{-}}\nu,&\nu\leq 0\\ \end{cases}.$ (15)
The pdf for the variable $\nu$ is a unit Gaussian, so for $R$ the pdf is a dimidiated (or split) Gaussian with two equal halves, as shown on the right of Figure 15. This is sometimes referred to as a bifurcated Gaussian, but that is certainly inaccurate. The dimidiated Gaussian for $R$ is the probability transform of the unit Gaussian for $\nu$, under the model that they are related by Equation (15)
@@
$p(x)=\begin{cases}{1\over{\sigma^{-}}\sqrt{2\pi}}e^{-{1\over 2}\left({x-M\over{\sigma^{-}}}\right)^{2}}&\text{for }x<M\\ {1\over{\sigma^{+}}\sqrt{2\pi}}e^{-{1\over 2}\left({x-M\over{\sigma^{+}}}\right)^{2}}&\text{for }x>M\\ \end{cases}.$ (16)
@@
$D=$ ${2\over 3S}\left[\sqrt{2\pi}\gamma-D^{3}({1\over\pi}-1)\right]$
@@
Even though the model can handle arbitrarily large asymmetries, in the sense of ${\sigma^{+}}-{\sigma^{-}}\over{\sigma^{+}}+{\sigma^{-}}$, there is a limit in terms of moments. If one of the two half-Gaussians has zero width, then
@@
so asymmetries, in the $\gamma\over{V^{3/2}}$ sense larger than this cannot be handled. But such cases are extreme.
A.1.1 Convolution of two dimidiated Gaussians